A recent 10-year study conducted by a research team at the Great Falls Medical School was conducted to assess how age, systolic blood pressure, and smoking relate to the risk of strokes. Assume that the following data are from a portion of this study. Risk is interpreted as the probability (times 100) that the patient will have a stroke over the next 10-year period. For the smoking variable, define a dummy variable with 1 indicating a smoker and 0 indicating a nonsmoker.

Risk Age Systolic Blood Pressure Smoker
12 57 152 no
24 67 163 no
13 58 155 no
56 86 177 yes
28 59 196 no
51 76 189 yes
18 56 155 yes
31 78 120 no
37 80 135 yes
15 78 98 no
22 71 152 no
36 70 173 yes
15 67 135 yes
48 77 209 yes
15 60 199 no
36 82 119 yes
8 66 166 no
34 80 125 yes
3 62 117 no
37 59 207 yes

A. What is the probability of a stroke over the next 10 years for Art Speen, a 65-year-old smoker who has a systolic blood pressure of 167? What action might the physician recommend for this patient?

B. An insurance company will only sell its Select policy to people for whom the probability of a stroke in the next ten years is less than .01. If a smoker with a systolic blood pressure of 230 applies for a Select policy, under what condition will the company sell him the policy if it adheres to this standard?

C. What other factors could be included in the model as independent variables? Choose the correct answer below.
(I) family history of stroke
   (ii) Weight/obesity
(iii) Gender
   (iv) All of the above
Answer #1

(a)

Predicted values for: Risk
Age Systolic Blood Pressure Smoker Predicted
65 167 1 29.021

Probability = 29.021/100 = 0.29021

(b) Risk = -91.7595 + 1.0767*Age + 0.2518*Systolic Blood Pressure + 8.7399*Smoker

0.01 = -91.7595 + 1.0767*Age + 0.2518*230 + 8.7399*1

Age = 23

The company sell him the policy if his age is 23.

(c)  (iv) All of the above